OpenStudy (anonymous):

During his workouts last week, Allen jogged the same amount of time each day for 6 days. Use the following information to determine the number of minutes Allen jogged each day: • During each workout, he jogged and he rode his bike. • He rode his bike for 45 minutes each day. • He worked out for a total of 360 minutes last week. How many minutes did Allen jog each day last week?

OpenStudy (anonymous):

Tell me if you understand this or not 45x+y=360 ( x = 6 ) y=360-264 y=96 One day he rode 96 minutes, the and the other six days he rode 45 minutes per day (a total of 264) If you want to email me anything EliLightE@Gmail.Com. I'll reply faster there.

OpenStudy (anonymous):

Ok Thanks X)